# Correlations versus noise in the NFT market

### v1

@@ III.2 Eigenvector components
where $Z_{1}(k)$ is the contribution to total variance associated with $\lambda_{1}$ ($k=1,...,T$), and diagonalizing the filtered matrix ${\bf C}^{\prime}$ constructed from the residual time series $\epsilon^{{}^{(i)}}(k)$ ($i=1,...,I$). The corresponding eigenvalue-eigenvector pairs for such a filtered matrix have an index decreased by unity $\lambda_{i}\rightarrow\lambda_{i-1}^{\prime}$, but for the sake of simplicity, we shall leave the index unchanged: $\lambda_{2}^{\prime},{\bf v}_{2}^{\prime}$. Fig. 8 shows the expansion coefficients $v_{2j}$ for the original matrix ${\bf C}$ without applying Eqs. (14) (top), for the filtered matrix ${\bf C}^{\prime}$ (middle), and for the detrended correlation matrix ${\bf C}^{\rho}(q,s)$ for $q=4$ and $s=14$ after filtering analogous to Eqs. (14) (bottom).
The first conclusion that can be drawn from the coefficients $v_{2j}$ is that Pudgy Penguins and the collection originated from it - Lil Pudgys, that were created by the same producer - TheIglooCompany, show strong correlations in terms of $c_{\Delta t}$. The same applies to ape-themed collections (Bored Ape Yacht Club, Bored Ape Kennel Club, and Mutant Ape Yacht Club) released by Yuga Labs. The related expansion coefficients are among the largest ones ($v_{2j}\in[0.2,0.4]$), together with Sappy Seals (top panel of Fig. 8). The close relations among collections released by TheIglooCompany become even more pronounced after removing the contribution of $\lambda_{1}$ (middle panel of Fig. 8), where their expansion coefficients $v_{2j}^{\prime}\approx 0.38]$ are much larger than for the other collections. The largest opposite-sign contribution to ${\bf v}_{2}^{\prime}$, which indicates a different behavior, is observed for Lives of Asuna collection (middle panel of Fig. 8). Other relations between collections can be inferred from ${\bf v}_{2}^{\rho}(q=4,s=14)$ (bottom panel of Fig. 8) for $q=4$, which amplifies large fluctuation contribution to the cross-correlations. In this case, $v_{2j}^{\rho}$ corresponding to Karafuru, 3Landers, and Killer GF collections that fall in the range $v_{2j}^{\rho}\in[0.38,0.4]$ and $v_{2j}^{\rho}$ corresponding to Goblintown that is equal to $v_{2j}^{\rho}\approx 0.25$ show significantly different values than the average $|v_{2j}^{\rho}|<0.2$ observed for the other collections. These less-liquid collections experienced a significant drop in capitalization on January 6, 2023, which was also accompanied by an increased number of transactions (inset in Fig 1(b)). The strongest opposite-sign contributions to ${\bf v}_{2}^{\rho}$ are made by Lil Pudgys and Sappy Seals ($v_{2j}^{\rho}\in[-0.15,-0.2]$).
@@ III.3 Minimal spanning trees (MST)
Figure 13: MST based on the $q$-dependent detrended correlation matrix ${\bf C}^{\rho}(q,s)$ for $q=4$ and $s=14$ for the capitalization increments $c_{\Delta t}$.
Figure 14: MST based on the $q$-dependent detrended correlation matrix ${\bf C}^{\rho}(q,s)$ for $q=4$ and $s=14$ for the number of transactions $N_{\Delta t}$.
@@
The $c_{\Delta t}$-based MST created from ${\bf C}^{\rho}(q,s)$ for $q=4$ and $s=14$ focuses on the correlation structure among large capitalization increments. Consistently with the fat-tailed distribution of the matrix elements depicted in Fig. 6(b), where there are few large elements and many small ones, the MST reveals a distributed structure with a few small clusters centered at Chain Runners ($\delta=11$), Robotos ($\delta=9$) and Cool Cats ($\delta=7$) – see Fig. 13. An evident property of the MSTs for $c_{\Delta t}$ is that the nodes corresponding to the collections with the largest capitalization (large circles), with the exception of the Ape collections, do not cluster together. Especially the collection with the largest capitalization – CryptoPunks – is peripheral in all the cases. This effect can be viewed as a contrasting one in respect to the other financial markets, where the largest stocks, currencies, and cryptocurrencies were identified as central nodes Kwapień and Drożdż (2012); Wątorek et al. (2021); Kwapień, Wątorek, and Drożdż (2021); Drożdż, Kwapień, and Wątorek (2023). For $N_{\Delta t}$, the corresponding MST is even more dispersed with the most connected nodes being Chain Runners ($\delta=7$) and Mutant Ape Yacht Club ($\delta=6$) – see Fig. 14. The cluster centered around the latter contains the collections with the largest number of transactions.
The decentralized structure of the MSTs presented in Figs. 9-14 can also be inferred from the node degree cumulative distributions shown in Fig. 15 where no node stands out as an outlier. In the case of the MSTs based on ${\bf C}$ (top panels), even a trace of a power-law dependence can be seen (straight lines in the double logarithmic scale plots). Such distributions of $\delta$ define the scale-free networks Ravasz and Barabási (2003) and were also reported for foreign currency exchange rates Górski, Drożdż, and Kwapień (2008) and cryptocurrencies Polovnikov, Kazakov, and Syntulsky (2020); Wątorek et al. (2021). No similar result can be seen for ${\bf C}^{\rho}(q,s)$ for $q=1$ (middle panels) and $q=4$ (bottom panels).
Figure 15: Cumulative distribution function $P(X\geq\delta)$ of the node degrees $\delta$ of MSTs created from the collection capitalization increments $c_{\Delta t}$ (left) and the number of transactions $N_{\Delta t}$ (right). The correlation matrices based on the Pearson coefficients ${\bf C}$ (top) and the $q$-dependent detrended correlation coefficients ${\bf C}^{\rho}(q,s)$ for $q=1$ (middle) and $q=4$ (bottom) were used.
@@ IV Summary
An additional insight into the cross-correlation structure of the NFT market has been given by a network approach that utilized minimal spanning trees. They reveal a largely decentralized structure of the respective MSTs that lack a dominant node, which remains in a striking contrast with the cryptocurrency market not to mention the traditional markets. Despite the fact that there exist large discrepancies in capitalization of the collections and the number of transactions, the price evolution of the largest, most frequently traded collections do not necessarily impact the respective characteristics of the less priced collections. This is yet another disparity between the NFT collections and the other financial assets, where the highly capitalized assets dominate the evolution of a market and tend to be coupled to each other.

### v2

@@ III.2 Eigenvector components
where $Z_{1}(k)$ is the contribution to total variance associated with $\lambda_{1}$ ($k=1,...,T$), and diagonalizing the filtered matrix ${\bf C}^{\prime}$ constructed from the residual time series $\epsilon^{{}^{(i)}}(k)$ ($i=1,...,I$). The corresponding eigenvalue-eigenvector pairs for such a filtered matrix have an index decreased by unity $\lambda_{i}\rightarrow\lambda_{i-1}^{\prime}$, but for the sake of simplicity, we shall leave the index unchanged: $\lambda_{2}^{\prime},{\bf v}_{2}^{\prime}$. Fig. 8 shows the expansion coefficients $v_{2j}$ for the original matrix ${\bf C}$ without applying Eqs. (14) (top), for the filtered matrix ${\bf C}^{\prime}$ (middle), and for the detrended correlation matrix ${\bf C}^{\rho}(q=4,s=14)$ after filtering analogous to Eqs. (14) (bottom).
The first conclusion that can be drawn from the coefficients $v_{2j}$ is that Pudgy Penguins and the collection originated from it - Lil Pudgys, that were created by the same producer - TheIglooCompany, show strong correlations in terms of $c_{\Delta t}$. The same applies to ape-themed collections (Bored Ape Yacht Club, Bored Ape Kennel Club, and Mutant Ape Yacht Club) released by Yuga Labs. The related expansion coefficients are among the largest ones ($v_{2j}\in[0.2,0.4]$), together with Sappy Seals (top panel of Fig. 8). The close relations among collections released by TheIglooCompany become even more pronounced after removing the contribution of $\lambda_{1}$ (middle panel of Fig. 8), where their expansion coefficients $v_{2j}^{\prime}\approx 0.38$ are much larger than for the other collections. The largest opposite-sign contribution to ${\bf v}_{2}^{\prime}$, which indicates a different behavior, is observed for Lives of Asuna collection (middle panel of Fig. 8). Other relations between collections can be inferred from ${\bf v}_{2}^{\rho}(q=4,s=14)$ (bottom panel of Fig. 8) for $q=4$, which amplifies large fluctuation contribution to the cross-correlations. In this case, $v_{2j}^{\rho}$ corresponding to Karafuru, 3Landers, and Killer GF collections that fall in the range $v_{2j}^{\rho}\in[0.38,0.4]$ and $v_{2j}^{\rho}$ corresponding to Goblintown that is equal to $v_{2j}^{\rho}\approx 0.25$ show significantly different values than the average $|v_{2j}^{\rho}|<0.2$ observed for the other collections. These less-liquid collections experienced a significant drop in capitalization on January 6, 2023, which was also accompanied by an increased number of transactions (inset in Fig 1(b)). The strongest opposite-sign contributions to ${\bf v}_{2}^{\rho}$ are made by Lil Pudgys and Sappy Seals ($v_{2j}^{\rho}\in[-0.15,-0.2]$).
@@ III.3 Minimal spanning trees (MST)
Figure 13: MST based on the $q$-dependent detrended correlation matrix ${\bf C}^{\rho}(q=4,s=14)$ for the capitalization increments $c_{\Delta t}$.
Figure 14: MST based on the $q$-dependent detrended correlation matrix ${\bf C}^{\rho}(q=4,s=14)$ for the number of transactions $N_{\Delta t}$.
@@
The $c_{\Delta t}$-based MST created from ${\bf C}^{\rho}(q=4,s=14)$ focuses on the correlation structure among large capitalization increments. Consistently with the fat-tailed distribution of the matrix elements depicted in Fig. 6(b), where there are few large elements and many small ones, the MST reveals a distributed structure with a few small clusters centered at Chain Runners ($\delta=11$), Robotos ($\delta=9$) and Cool Cats ($\delta=7$) – see Fig. 13. An evident property of the MSTs for $c_{\Delta t}$ is that the nodes corresponding to the collections with the largest capitalization (large circles), with the exception of the Ape collections, do not cluster together. Especially the collection with the largest capitalization – CryptoPunks – is peripheral in all the cases. This effect can be viewed as a contrasting one in respect to the other financial markets, where the largest stocks, currencies, and cryptocurrencies were identified as central nodes Kwapień and Drożdż (2012); Wątorek et al. (2021); Kwapień, Wątorek, and Drożdż (2021); Drożdż, Kwapień, and Wątorek (2023). For $N_{\Delta t}$, the corresponding MST is even more dispersed with the most connected nodes being Chain Runners ($\delta=7$) and Mutant Ape Yacht Club ($\delta=6$) – see Fig. 14. The cluster centered around the latter contains the collections with the largest number of transactions.
The decentralized structure of the MSTs presented in Figs. 9-14 can also be inferred from the node degree cumulative distributions shown in Fig. 15 where no node stands out as an outlier. In the case of the MSTs based on ${\bf C}$ (top panels), even a trace of a power-law dependence can be seen (straight lines in the double logarithmic scale plots). Such distributions of $\delta$ define the scale-free networks Ravasz and Barabási (2003) and were also reported for foreign currency exchange rates Górski, Drożdż, and Kwapień (2008) and cryptocurrencies Polovnikov, Kazakov, and Syntulsky (2020); Wątorek et al. (2021). No similar result can be seen for ${\bf C}^{\rho}(q=1,s)$ (middle panels) and ${\bf C}^{\rho}(q=4,s)$ (bottom panels).
Figure 15: Complementary cumulative distribution function $P(X\geq\delta)$ of the node degrees $\delta$ of MSTs created from the collection capitalization increments $c_{\Delta t}$ (left) and the number of transactions $N_{\Delta t}$ (right). The correlation matrices based on the Pearson coefficients ${\bf C}$ (top) and the $q$-dependent detrended correlation coefficients ${\bf C}^{\rho}(q=1,s)$ (middle) and ${\bf C}^{\rho}(q=4,s)$ (bottom) were used.
@@ IV Summary
An additional insight into the cross-correlation structure of the NFT market has been given by a network approach that utilized minimal spanning trees. They reveal a largely decentralized structure of the respective MSTs that lack a dominant node, which remains in a striking contrast with the cryptocurrency market not to mention the traditional markets. Despite the fact that there exist large discrepancies in capitalization of the collections and the number of transactions, the price evolution of the largest, most frequently traded collections are not necessarily associated with the respective characteristics of the less priced collections. This is yet another disparity between the NFT collections and the other financial assets, where the highly capitalized assets dominate the evolution of a market and tend to be coupled to each other.